OpenStudy (anonymous):

Solve: x2 - 20x + 36 = 0 (Points : 1) {6, -6} {2, 18} {-4, -9} no solution

OpenStudy (campbell_st):

just substitute each x value e.g 6^2 - 20(6) + 36 = -48 -6^2 - 20(-6) + 36 = 192